Two spherical objects have masses of 1.5 x 105 kg and 8.5 x 102 kg. Their centers are separated by a distance of 2500 m. Find the gravitational attraction between them.


1
Expert's answer
2022-02-11T08:47:05-0500

Given:

"m_1=1.5 *10^5\:\rm kg"

"m_2=8.5 *10^2\:\rm kg"

"r=2500\:\rm m"


The law of universal gravitation says

"F=G\frac{m_1m_2}{r^2}""F=6.67*10^{-11}\frac{1.5*10^5*8.5*10^2}{2500^2}=1.4*10^{-9}\:\rm N"